OpenStudy (anonymous):

can anyone explain to me how to do this: 2x^2 - 54 = 284

OpenStudy (anonymous):

this is what i got but i dont think it's right; x^2 - 54 = 142 +54 +54 -------------- x^2 = 196 Square Root of 196 = 14 ?

OpenStudy (campbell_st):

add 54 to both sides 2x^2 = 338 divide by 2 x^2 = 169 take the square root which will give a positive and negative solution \[x=\sqrt{169} = \pm 13\]

OpenStudy (anonymous):

ooh, i divide by 2 first.

OpenStudy (anonymous):

thanks!! :)

OpenStudy (campbell_st):

54/2 = 27

OpenStudy (anonymous):

and can you check 2 more for me? :)

OpenStudy (campbell_st):

yep

OpenStudy (anonymous):

7n^2 = 63 Divided by 6 = n^2 = 9 n = (3, -3)

OpenStudy (campbell_st):

divide by 7 n^2 = 9 \[n = \sqrt{9}= \pm 3\]

OpenStudy (anonymous):

okay thank you! one more problem.

OpenStudy (anonymous):

can you explain to me how to solve this kind of problem?